### v2

@@ 5.1 Agent Design
We implement a simple web agent with chain-of-thought prompting (Wei et al., 2022b), and we evaluate its performance across two axes: (1) the underlying LLM, and (2) the use of BrowserGym features.
Observation space:  The observation space is composed of the goal, the DOM formatted in HTML, the accessibility tree and the errors of the previous action if any. We also provide the history of actions since the beginning of the episode. To study the effect of the different components, we use flags to activate or deactivate certain features such as use_error_logs, use_html, use_AXTree. We have the option to augment the HTML and accessibility tree with the coordinates of each object on the page. This is achieved using extract_center or extract_box to extract the centre coordinates or bounding boxes respectively. Similarly, we can extract whether objects are visible or hidden using extract_visible_tag.
Action space:  Agents can be designed to use single actions only or multi-actions. When action_space=bid, only actions to interact at the object level are permitted. When action_space=bid+coord, we augment the set of actions to interact with 2D coordinates $x,y$. This is useful in MiniWoB, where certain tasks require clicking at a certain position on a SVG figure.
@@
Zero-shot examples:  In the prompt, we provide a single generic example of how the chain of thought and the action should be formatted. This contrasts with other methods (Kim et al., 2023) where task-specific few-shot examples are provided, yet aligns with our objective of developing zero-shot agents able to solve a large range of new tasks.
@@
Parse and retry:  Once the LLM provides an answer, we have a parsing loop that can re-prompt the agent up to 4 times to make it aware of the parsing mistake. This can save the agent from making basic mistakes and is mainly useful for less capable LLMs such as GPT-3.5 and CodeLlama. Once parsed, the action is executed via BrowserGym, which moves to the next step.
@@
Language models:  Our study distinguishes between closed- and open-source LLMs. For the closed-source segment, we employed GPT-3.5 (gpt-3.5-turbo-1106, 16K context) and GPT-4 (OpenAI, 2023) (gpt-4-1106-preview, 128K context), utilizing OpenAI’s API for model instantiation. In the realm of open-source LLMs, we sought a model that 1) understands code and HTML, 2) can manage our lengthy prompts with a substantial context size, and 3) is instruction-finetuned. Our choice fell on CodeLLAMA-34b-instruct, a specialized version of LLAMA2 (Touvron et al., 2023) that is finetuned for code, extended context, and instruction-following capabilities. This model was deployed using Hugging Face’s Text Generation Inference library on 8 A100 GPUs, setting the stage for a maximum single-batch size of 12,000 tokens. We also explored the effect of using GPT-4 vision by providing the screenshot of the page but only observed marginal improvement on MiniWoB.
Truncating prompt:  We use a maximum of 15,000 tokens for our prompt when using GPT-3.5, 11,000 for CodeLlama and 40,000 with GPT-4. When the prompt is too large, we progressively truncate the DOM and accessibility tree until they fit the maximum allowed.
@@ 5.2 Experimental Protocol
Standard Error:  To be able to run a range of experiments under a fixed budget, we limit the number of seeds to 10 per task. After averaging results over each benchmark (or subset of it), we usually observe a sufficiently low standard error to draw the needed conclusions. We use stratified bootstrap [^4] We sample with replacement for each task independently and we average across the tasks and seeds. to obtain 1,000 samples of the mean and report the average and standard deviation of these means as success rate and standard error.
@@
Max step:  For all tasks, we allow a maximum of 10 steps per episode. This ensures that a low-performing agent will not wander around for too long if it’s incapable of solving the task; 10 steps are considered sufficient for all MiniWoB tasks. On the other hand, some WorkArena tasks can require a single-action agent more than 25 steps to accomplish, although 10 are enough in multi-action mode.
Model Selection:  To select the best model from the ablation study, we use 3 seeds out of the 10 and keep the remaining 7 for the bootstrap statistics.
5.3 Baselines on WorkArena
Tab. 1 reports the zero-shot performance of these agents on three benchmarks: MiniWoB (Liu et al., 2018), WebArena (Zhou et al., 2023) and our proposed WorkArena. Performance is measured in terms of success rate. We emphasize our key findings below.
Table 1: Success rate±$\pm$±Standard error (SR $\pm$±SE) of all agents on MiniWoB, WorkArena and WebArena. Bolded numbers represent the average success rate over the entire corresponding benchmark.
[…]
Task Category GPT-4 GPT-3.5 CodeLlama
SR % $\pm$±SE SR % $\pm$±SE SR % $\pm$±SE
WorkArena 54.8 $\pm$2.1 18.6 $\pm$2.2 0 $\pm$0
   Form 58.0 $\pm$4.8 16.0 $\pm$4.0 0 $\pm$0
   Knowledge 50.0 $\pm$14.8 0.0 $\pm$4.3 0 $\pm$0
   List-filter 0.0 $\pm$1.7 0.0 $\pm$2.0 0 $\pm$0
   List-sort 58.3 $\pm$5.5 38.3 $\pm$6.1 0 $\pm$0
   Menu 95.0 $\pm$4.8 25.0 $\pm$8.7 0 $\pm$0
   Service catalog 78.9 $\pm$3.7 20.0 $\pm$3.3 0 $\pm$0
MiniWoB (125 tasks) 71.7 $\pm$1.0 43.6 $\pm$1.0 25.5 $\pm$1.3
   WebGum Subset (56 tasks) 87.6 $\pm$1.2 59.8 $\pm$1.7 32.4 $\pm$2.1
WebArena (812 tasks) 25.4 $\pm$0.6 13.1 $\pm$0.6 0 $\pm$0
   Content-and-config (411 tasks) 28.7 $\pm$0.9 21.9 $\pm$0.9 0 $\pm$0
   Information-seeking (325 tasks) 23.1 $\pm$1.0 4.0 $\pm$1.0 0 $\pm$0
   Navigation (76 tasks) 17.1 $\pm$2.1 3.9 $\pm$2.0 0 $\pm$0
Performance on MiniWoB:  The GPT-4-based agent demonstrates notably high performance on this benchmark, achieving significantly greater zero-shot success compared to other agents. The outcomes on the full benchmark and the WebGum subset surpass those of prior studies on zero-shot performance like Assouel et al. (2023); Zeng et al. (2023), underscoring the effectiveness of our agent design.
WorkArena Presents a Greater Challenge:  Consistent with our expectations, our newly proposed benchmark, WorkArena, proves to be significantly more difficult than MiniWoB, primarily due to its incorporation of complex user interfaces from real-world software environments. Consequently, all agents exhibit low performance levels, with no agent achieving success in one specific task. Notably, our top-performing agent was only able to exceed a 50% success rate in 2 out of the 6 categories, highlighting the benchmark’s heightened challenge.
CodeLLAMA’s Shortcomings:  Our analysis sheds light on why CodeLLAMA, despite its reasonable zero-shot performance on MiniWoB, fails to accomplish any task within WorkArena. We categorize and summarize the primary failure modes encountered, with details in § C. Initially, CodeLLAMA frequently replicates actions from prompt examples, which are unsuitable for a given context (§ C.1). Additionally, it misapplies both high-level (§ C.2) and Python APIs (§ C.3), and struggles with the intricacies of handling iFrames in Playwright (§ C.3). On the rare occasions that CodeLLAMA does execute relevant actions and navigates a WorkArena instance, it nevertheless fails to meet its objectives. These observations underscore a significant disparity between leading open-source LLMs and their closed counterparts further highlighting the rigorous challenge posed by our benchmark.
@@ 5.3 Baselines on WorkArena
GPT-4’s Superiority:  The data unequivocally demonstrate GPT-4’s dominance over GPT-3.5 and CodeLlama within WorkArena. In the relatively simpler task category of menu navigation, GPT-4 achieves 95% success, in stark contrast to GPT-3.5’s 25% and CodeLlama’s 0%. The performance disparity between the LLMs is significantly more pronounced here than in MiniWoB and WebArena, aligning with findings on the emergent properties of AI systems (Wei et al., 2022a). As the complexity of tasks increases, the necessity for a more advanced LLM to achieve any score becomes apparent, with noticeable improvements in performance correlating with the enhancement of capabilities.
5.4 Challenges of Large Context in WorkArena
WorkArena underscores the challenge of managing real-world websites, notably dealing with extensive observations stemming from sizable HTML files, as illustrated in Fig. 4. This poses a particular challenge for Large Language Models (LLMs) that require encoding this vast context to make accurate decisions. Moreover, current LLMs rely on the transformer architecture, which incurs a quadratic cost in attention computation. For instance, a typical 100,000 token HTML file in WorkArena necessitates a transformer to compute 10 billion ($100,000^{2}$) attention values per layer to generate the first token, significantly impacting computational efficiency.
Fortunately, BrowserGym introduces an accessibility representation of websites (Zhou et al., 2023), significantly diminishing the context size required for processing. Additionally, through the implementation of visible element tagging, it is possible to prune elements that are not visible to the user, thereby further reducing the context size. This approach not only mitigates the computational challenges posed by large HTML files but also aligns more closely with the actual user experience by focusing on elements that are relevant and visible, enhancing the efficiency and relevance of the model’s actions within the environment.
Figure 4: Comparative analysis of observation modality sizes: (left) DOM elements of the page and (right) its accessibility tree, across MiniWoB and WorkArena. In WorkArena, the data is further delineated by task category. The grey regions indicate the augmented observation size when incorporating the comprehensive suite of features provided by BrowserGym.
@@ 5.4 Challenges of Large Context in WorkArena
Table 2: Progressive feature analysis on MiniWoB. Success rate±$\pm$±Standard error (SR $\pm$±SE) of all configurations. Each row adds a new feature to the previous configuration.
Configuration GPT-4 GPT-3.5
@@
Most basic configuration 65.1 $\pm$0.9 29.7 $\pm$0.9
$\hookrightarrow$ use_error_logs=True 66.6 $\pm$1.0 32.2 $\pm$1.0
$\hookrightarrow$ use_ax_tree=True 66.1 $\pm$1.0 38.2 $\pm$1.0
$\hookrightarrow$ multi_actions=True 67.9 $\pm$1.0 42.0 $\pm$1.1
$\hookrightarrow$ extract_coords=center 70.4 $\pm$0.8 43.6 $\pm$1.0
$\hookrightarrow$ action_space=bid+coord 68.4 $\pm$1.0 41.6 $\pm$1.1
$\hookrightarrow$ extract_coords=box 71.7 $\pm$1.0 39.1 $\pm$1.1
$\hookrightarrow$ extract_visible_tag=True 66.9 $\pm$1.1 39.8 $\pm$1.1
Table 3: Progressive feature analysis on WorkArena. Success rate±$\pm$±Standard error (SR $\pm$±SE) of all configurations. Each row adds a new feature to the previous configuration.
[…]
Configuration GPT-4 GPT-3.5
@@
Most basic configuration 50.7 $\pm$2.0 13.8 $\pm$1.7
$\hookrightarrow$ use_error_logs=True 54.8 $\pm$2.0 14.5 $\pm$1.6
$\hookrightarrow$ multi_actions=True 50.2 $\pm$2.0 18.6 $\pm$2.2
$\hookrightarrow$ extract_coords=center 51.2 $\pm$2.3 14.5 $\pm$1.7
$\hookrightarrow$ action_space=bid+coord 51.2 $\pm$1.9 13.8 $\pm$1.7
$\hookrightarrow$ extract_coords=box 50.7 $\pm$2.4 13.1 $\pm$1.8
$\hookrightarrow$ extract_visible_tag=True 50.3 $\pm$2.0 13.1 $\pm$1.8
5.5 Effect of BrowserGym Features
Our proposed BrowswerGym environment equips web agents with rich observation and action spaces. Results reported in Tabs. 2 and 3 show the effect of combinations of such features on performance on MiniWoB and WorkArena, respectively. Here, the most basic configuration is where all the features are turned off and each row corresponds to increasingly adding more features (cf.§ 5.1).
BrowserGym improves performance:  Tabs. 2 and 3 shows that adding observation- or action-space features often helps performance. While not always the case, we can observe a gap of more than 10% between the most basic configuration and the best one using GPT-3.5 on MiniWoB. Moreover, our GPT4 agents obtains state-of-the-art performance on WebArena with 25.4% success rate.
2D features do not help for WorkArena:  Several MiniWoB tasks require 2D understanding of the UI and interaction with $x,y$ coordinates; we observe a notable performance gain when introducing 2D features in Tab. 2. On the other hand, on WorkArena the tasks seem best solved using bid only.

### v5

@@ 5.1 Agent Design
We implement a simple web agent with chain-of-thought prompting (Wei et al., 2022b), and evaluate its performance across two axes: (1) the underlying LLM, and (2) the use of BrowserGym features. To study the effect of different design choices in our agents, we use flags to activate or deactivate certain features, such as use_thinking for chain-of-thought.
Observation space:  Our observation space is composed of the goal, the current page’s HTML and/or AXTree, [^6] On WebArena and WorkArena we only use AXTrees because HTML is prohibitively large. On MiniWoB we use both AXTree and HTML as it consistently gave the best performance. the currently focused element, and the error from the previous action if any. We activate / deactivate these features with flags use_last_error and use_focused_element, and we decide whether or not to augment each element with extra information using flags coords=center or coords=box for screen coordinates, and extract_visible_tag and extract_clickable_tag for whether elements are visible and/or clickable. Last, the flag filter_visible_only allow us to only include elements that are visible on the screen.
Action space:  We leverage the flexibility of BrowserGym with a multi_action flag, which either allows agents to execute multiple actions per step (e.g., click(12) click(52)) or not. The action_set=bid flag only permits primitives that use element identifiers, while action_set=bid+coord also includes primitives operating with 2D coordinates (x,y). This is useful in MiniWoB, where certain tasks require clicking at specific positions in SVG images. Finally, we also experiment with the textual description of the action space which can either include a short or long description for each primitive (long_description), as well as examples of valid function calls or not (individual_examples).
History:  To extend the horizon window of our agents, we experiment with use_action_history which includes in the prompt the history of actions since the beginning of the episode, use_error_history which includes all previous error messages, and use_think_history which re-injects the previous chain-of-though outputs into the prompt, thus creating an artificial memory that gives agents a chance to recall their previous thoughs.
@@
Zero-shot examples:  In the prompt, we provide a single generic example of how the chain of thought and action outputs should be formatted. This contrasts with other methods (Kim et al., 2023) where task-specific few-shot examples are provided, yet aligns with our objective of developing zero-shot agents able to solve a large range of new tasks.
@@
Parse and retry:  Once the LLM provides an answer, we have a parsing loop that can re-prompt the agent up to 4 times to make it aware of a parsing mistake. This can save the agent from making basic mistakes and is mainly useful for less capable LLMs such as GPT-3.5. Once parsed, the action is executed via BrowserGym, which moves to the next step.
@@
Language models:  Our study distinguishes between closed- and open-source LLMs. For the closed-source segment, we evaluate GPT-3.5 (gpt-3.5-turbo-1106, 16K context) and GPT-4o (OpenAI, 2023) (gpt-4o-2024-05-13, 128K context), through OpenAI’s API. In the realm of open-source LLMs, we sought a model that 1) understands code and HTML, 2) can manage a substantial context size, and 3) is instruction-finetuned. Our choice fell on Llama3-70b (Meta, 2024) (meta-llama-3-70B-instruct, 8K context), a recently released model that exhibits performances close to GPT-4. This model was deployed using Hugging Face’s Text Generation Inference (TGI) library on 4 A100 GPUs. We also explore the effect of providing the screenshot of the page using GPT-4o vision and Set-of-Mark (Yang et al., 2023) as proposed in WebVoyager (He et al., 2024).
Prompt truncation:  We use a maximum prompt length of 40K tokens for GPT-4o, 15K for GPT-3.5 and 8K for Llama3. When the prompt is too large (Figure 4), we progressively truncate the HTML and AXTree from the end until it fits the maximum allowed number of tokens.
Table 2: Success rate±$\pm$±Standard error (SR $\pm$±SE) of all agents on MiniWoB, WorkArena, and WebArena. Bolded numbers represent the average success rate over the entire corresponding benchmark.
Task Category GPT-4o GPT-4o-V GPT-3.5 Llama3
SR % $\pm$±SE SR % $\pm$±SE SR % $\pm$±SE SR % $\pm$±SE
WorkArena (33 tasks) 42.7 $\pm$1.5 41.8 $\pm$1.7 6.1 $\pm$1.3 17.9 $\pm$1.5
   Dashboard (4) 62.5 $\pm$6.8 72.5 $\pm$6.0 20.0 $\pm$4.8 37.5 $\pm$6.0
   Form (5) 40.0 $\pm$5.9 34.0 $\pm$4.8 2.0 $\pm$2.5 32.0 $\pm$4.6
   Knowledge (1) 80.0 $\pm$12.2 70.0 $\pm$13.9 0.0 $\pm$4.3 30.0 $\pm$12.3
   List-filter (6) 0.0 $\pm$1.6 0.0 $\pm$1.7 0.0 $\pm$1.6 0.0 $\pm$1.8
   List-sort (6) 10.0 $\pm$3.8 13.3 $\pm$4.0 8.3 $\pm$3.7 1.7 $\pm$2.5
   Menu (2) 60.0 $\pm$8.0 90.0 $\pm$6.0 5.0 $\pm$4.7 0.0 $\pm$2.9
   Service catalog (9) 77.8 $\pm$3.2 65.6 $\pm$3.6 5.6 $\pm$2.3 26.7 $\pm$3.4
MiniWoB (125 tasks) 66.1 $\pm$1.0 67.7 $\pm$1.0 38.9 $\pm$1.1 62.6 $\pm$0.6
   WebGum Subset (56) 82.9 $\pm$1.5 83.2 $\pm$1.5 53.6 $\pm$1.4 80.5 $\pm$1.0
WebArena (812 tasks) 23.5 $\pm$0.7 24.0 $\pm$0.6 6.7 $\pm$0.6 11.0 $\pm$0.6
   Content-and-config (411) 25.8 $\pm$1.0 26.8 $\pm$0.9 8.8 $\pm$0.8 12.7 $\pm$0.9
   Information-seeking (325) 22.5 $\pm$1.0 22.5 $\pm$0.9 4.3 $\pm$0.9 9.8 $\pm$1.1
   Navigation (76) 15.8 $\pm$2.2 15.8 $\pm$1.8 5.3 $\pm$1.9 6.6 $\pm$1.9
@@ 5.2 Experimental Protocol
Standard Error:  To be able to run a range of experiments under a fixed budget, we limit the number of seeds to 10 per task for MiniWoB and WorkArena, and 1 per task for WebArena. After averaging results over each benchmark (or subset of it), we usually observe a sufficiently low standard error to draw the needed conclusions. We use stratified bootstrap to obtain 1,000 samples of the mean and report the average and standard deviation of these means as success rate and standard error.
@@
Max step:  In all tasks we give our agents a maximum of 15 steps per episode, as in He et al. (2024). This ensures a low-performing agent will not wander for too long if they are incapable of solving the task. Note that 15 steps is considered sufficient for MiniWoB, but on WorkArena some tasks might require more than 15 steps to complete unless the agent runs in multi-action mode (see § A.1 Table 6 for an upper bound).
Model selection:  We evaluate all LLMs using a common agent code-base that we tune by activating / deactivating flags (see § 5.1). For each LLM we find the best agent configuration via random search on MiniWoB and WorkArena. We then fix the final configuration (one per LLM), and re-run a full evaluation on all benchmarks with a different seed. [^7] WebArena being deterministic, we don’t use it for tuning. Final configurations are reported in Table 1.
5.3 Results
We report the final performance of our agents on MiniWoB (Liu et al., 2018), WebArena (Zhou et al., 2023) and WorkArena in Table 2. We emphasize our key findings below.
@@ 5.3 Results
GPT-4o’s Superiority:  The data unequivocally demonstrate GPT-4o’s dominance over GPT-3.5 and Llama3, on all benchmarks. The performance gap is particularly striking on WorkArena, where GPT-4o achieves 43% success, in stark contrast to GPT-3.5’s 6.1% and Llama3’s 17.9%. The performance disparity between LLMs is significantly more pronounced in WorkArena and WebArena than in MiniWoB, aligning with findings on the emergent properties of AI systems (Wei et al., 2022a). As the complexity of task increases, the necessity for a more advanced LLM to achieve any score becomes apparent, with noticeable improvements in performance correlating with the enhancement of capabilities.
WorkArena poses a great challenge:  Consistent with our expectations, our newly proposed benchmark, WorkArena, proves to be a significant challenge for current LLMs. While the tasks remain simple at a high-level (navigating menus, filling forms) and mostly require general knowledge, the difficulty arises primarily from the use of complex user interfaces from real-world software environments, resulting in long contexts and non-trivial interactions. Consequently, all agents exhibit low performance levels, with no agent achieving 100% success in one specific task. Notably, the list-based tasks which require interacting with a non-standard HTML widget (§ A.2 Figure 5) prove to be the most challenging ones with a clear 0% success rate for all LLMs, while being fairly simple for a human to complete.
State-of-the-art on MiniWoB and WebArena:  Our GPT-4o agent demonstrates notably high performance on MiniWoB, achieving significantly greater success compared to other agents. The outcomes on the full benchmark and the WebGum subset (82.9%) surpass those of prior studies on zero-shot web agents (Assouel et al., 2023; Zeng et al., 2023), underscoring the effectiveness of both our agent design and the features offered by BrowserGym. On WebArena, while the score obtained by our GPT-4o agent is only 23.5%, it constitutes the best zero-shot performance reported so far, way above the 14.4% succes rate obtained using GPT-4 in the original paper(Zhou et al., 2023).
Open-source LLMs:  While our Llama3 agent still performs poorer than a closed-source GPT-4o, its performance is remarkably high compared to a GPT-3.5 agent despite its shorter context window (8K vs 16K). Most notably, in preliminary experiments with a Llama2 model we could not obtain any success (flat 0%) on both WebArena and WorkArena, which makes us hopeful for the future of open-source LLMs. These results also indicate that web agent benchmarks, and WorkArena in particular, offer a great tool for evaluating emerging capabilities in LLMs, with larger performance gaps and more room for improvement compared to regular benchmarks (Touvron et al., 2023).
@@
Table 3: Ablation study for GPT-4o on MiniWoB and WorkArena. Success rate±$\pm$±Standard error (SR $\pm$±SE) of all configurations. Each row modifies the initial configuration.
Configuration MiniWoB WorkArena
@@
Initial configuration 68.2 $\pm$1.0 45.5 $\pm$2.2
+multi_actions 68.5 $\pm$1.0 40.6 $\pm$2.0
+coords=box,action_set=bid+coord 72.6 $\pm$1.0 41.2 $\pm$1.8
+use_think_history 66.7 $\pm$0.9 42.4 $\pm$2.3
+use_error_history 67.2 $\pm$0.9 43.6 $\pm$2.1
-extract_visible_tag 68.8 $\pm$1.0 43.0 $\pm$2.2
Table 4: Ablation study for GPT-3.5 on MiniWoB and WorkArena. Success rate±$\pm$±Standard error (SR $\pm$±SE) of all configurations. Each row modifies the initial configuration.
Configuration MiniWoB WorkArena
@@
Initial configuration 41.3 $\pm$1.1 8.5 $\pm$1.3
-use_thinking 30.2 $\pm$1.0 6.1 $\pm$1.2
-use_action_history 34.1 $\pm$1.1 5.2 $\pm$1.3
+use_think_history 36.2 $\pm$1.2 3.0 $\pm$1.0
+use_error_history 37.9 $\pm$1.0 7.0 $\pm$1.4
+multi_actions 41.4 $\pm$1.1 9.4 $\pm$1.5
+only_visible_elements 38.1 $\pm$1.0 5.8 $\pm$1.3
+long_description 39.5 $\pm$1.0 6.4 $\pm$1.2
-individual_examples 34.4 $\pm$1.0 8.2 $\pm$1.3
+coords=center, action_set=bid+coord 40.8 $\pm$1.1 4.5 $\pm$1.1
+coords=box, action_set=bid+coord 40.5 $\pm$1.2 6.1 $\pm$1.4
+use_focused_element 40.3 $\pm$1.1 9.4 $\pm$1.5
-extract_visible_tag 37.8 $\pm$1.0 6.7 $\pm$1.3
+extract_clickable_tag 37.9 $\pm$1.1 7.6 $\pm$1.2
Table 5: Ablation study for Llama3 on MiniWoB and WorkArena. Success rate±$\pm$±Standard error (SR $\pm$±SE) of all configurations. Each row modifies the initial configuration.
Configuration MiniWoB WorkArena
SR % $\pm$±SE SR % $\pm$±SE
Initial configuration 59.8 $\pm$1.0 20.0 $\pm$2.3
-use_thinking 48.6 $\pm$1.1 8.5 $\pm$1.7
-use_action_history 49.9 $\pm$1.1 8.5 $\pm$1.7
-use_think_history 52.2 $\pm$1.1 18.8 $\pm$2.1
+multi_actions 63.0 $\pm$1.0 17.6 $\pm$2.1
+long_description 55.5 $\pm$1.0 17.6 $\pm$2.0
+use_last_error 60.2 $\pm$1.1 19.4 $\pm$1.8
-extract_visible_tag 61.3 $\pm$0.9 15.8 $\pm$2.0
5.4 Ablation Study
We report in Tables 3, 4 and 5 an ablation study of each LLM on MiniWoB and WorkArena, for the most important agent flags. [^8] The ablation study was done with a different seed, hence numbers differ slightly from those in Table 2. We present our findings below.
Chain-of-though is crucial:  In all of our 3 agents, asking the LLM to directly produce the next best action without chain-of-though is very detrimental to performance. This is clearly seen in Table 5, with a drop of 10 points for Llama3 on both MiniWoB and WorkArena. This conclusion holds for all three LLMs very clearly.